OpenStudy (anonymous):

during the first part of a trip, a canoeist travels 88miles at a certain speed. the canoeist travels 10 miles on the second part of the trip at a speed 5mph slower. total time for the trip is 4hrs. what was the speed on both parts of the trip?

OpenStudy (anonymous):

Let r be the speed for the first 88 miles. Solve the following equation for r:\[\frac{88}{r}+\frac{10}{r-5}=4\]\[ r=25.12 \text{ mph}, r-5=20.12\text{ mph}\]